The physician has ordered IV heparin sodium 25,000 units in 1000mL of 0.9% Sodium Chloride to infuse at a rate of 1000 units of heparin per hour. What will your flow rate be to infuse 1000 units/hour?
Answer #1

ANSWER.

* Calculation of flow rate of heparin sodium.

          Formula = Desired dose / Dose in hand * Quantity.

                 Desired dose(D) = 1000 units/hr.

                 Dose in hand(H) = 25000 units.

                        Quantity (Q) = 1000 ml.

        As per formula

             Flow rate to infuse 1000 unit/hr is = D/H * Q

                                = 1000 / 25000 * 1000

                            = 0.04 * 1000

                          = 40 ml/hr.